Documenta Mathematica

Band 22, 2017

E. Bayer-Fluckiger and R. Parimala Cohomological Invariants for G-Galois Algebras

and Self-Dual Normal Bases  1–24

Matthias Nagel, Mark Powell Concordance Invariance of

Levine-Tristram Signatures of Links  25–43

Mikhail Bondarko and Gonçalo Tabuada Picard Groups, Weight Structures,

and (noncommutative) Mixed Motives  45–66

Lidia Angeleri Hügel and Dirk Kussin Large Tilting Sheaves over Weighted

Noncommutative Regular Projective Curves  67–134

A.D. Măcinic, Ş. Papadima, C.R. Popescu Modular Equalities

for Complex Reflection Arrangements  135–150

Finnur Lárusson and Tuyen Trung Truong Algebraic Subellipticity and Dominability

of Blow-Ups of Affine Spaces  151–163

iv